If the real interest rate is 7.3% and the inflation rate is 4.2%, what is the nominal interest rate? Enter you answer as a percentage. Do not enter the percentage sign into your answer Enter your response below (rounded to 2 decimal places)

Please show your work
Answer #1

(1 + NOMINAL RATE) = (1+ REAL RATE)*(1+ INFLATION RATE)

(1 + NOMINAL RATE) = (1+ 0.073)*(1+ 0.042) = 1.118066

NOMINAL RATE = 1.118066 - 1 = 0.118066 = 11.8066% = 11.81%

ANSWER : 11.81% (THUMBS UP PLEASE)